OpenStudy (anonymous):

What is the solution of 2^2x = 8x + ^1?

OpenStudy (anonymous):

Option A: -4 Option B: -3 Option C: -1 Option D: 1/2

OpenStudy (anonymous):

is that how its actually written?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just ^1 not x^1

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Which is why I am confused.

OpenStudy (anonymous):

lol im with you on that

OpenStudy (anonymous):

If you can help answer though it would be fantasic

OpenStudy (anonymous):

can you screen shot the question?

OpenStudy (anonymous):

sure give me a moment

OpenStudy (anonymous):

there you go :)

OpenStudy (anonymous):

i feel like its 3 but i don't want you to get it wrong

OpenStudy (xapproachesinfinity):

this is what you wrote dear \[2^{2x}=8x+ ^1\]

OpenStudy (anonymous):

Oh my gosh I didnt meant too oops

OpenStudy (xapproachesinfinity):

be careful when you write thing you want to make sure it's written well

OpenStudy (xapproachesinfinity):

no problem back to your problem

OpenStudy (xapproachesinfinity):

what ideas have tried so far

OpenStudy (xapproachesinfinity):

any work?

OpenStudy (xapproachesinfinity):

i want you to look at the problem and tell me what can you do show me some work in other words

OpenStudy (anonymous):

I am not sure how to approach the problem. We only did a few in class and I didn't understand it then either.

OpenStudy (xapproachesinfinity):

well you should probably start from the basics you can't attack a problem with nothing in mind you have to have some baggage with you

OpenStudy (anonymous):

sorry i meant -3

OpenStudy (xapproachesinfinity):

let's pick it from some algebra knowledge \[((a)^n)^m\]

OpenStudy (xapproachesinfinity):

do you know hod to deal with this ?

OpenStudy (anonymous):

I probably should, but I dont. :/

OpenStudy (xapproachesinfinity):

well true you have done algebra so this should be obvious to you it gonna be \[a^{n\times m}\]

OpenStudy (xapproachesinfinity):

now you do want to keep that one in mind because algebra is key to all the stuff in math

OpenStudy (xapproachesinfinity):

in our problem here we want to go back words

OpenStudy (xapproachesinfinity):

we are doing this to the right hand side okay

OpenStudy (anonymous):

Okay! :)

OpenStudy (xapproachesinfinity):

\[2^{2x}=8^{x+1} \Longrightarrow 2^{2x}=(2^3)^{x+1}\]

OpenStudy (xapproachesinfinity):

i took 8 and made is 2^3 but 8 is raised to power x+1 so 2^3 is raised to x+1 as well

OpenStudy (xapproachesinfinity):

good so far

OpenStudy (anonymous):

Well, why did you make the 8 to a 2^3

OpenStudy (xapproachesinfinity):

well you have to know in the back of you mind that 8=2x2x2 which the same as 2^3 so i used that above

OpenStudy (anonymous):

(I am trying to understand. Im sorry.)

OpenStudy (anonymous):

Oh okay

OpenStudy (xapproachesinfinity):

yeah i just explained never get afraid of asking :)

OpenStudy (xapproachesinfinity):

so we agree that 2^3=8

OpenStudy (anonymous):

Yes okay, I understand

OpenStudy (xapproachesinfinity):

good! now go back to what i said earlier about some algebra skills \[(a^n)^m=a^{n\times m}\] we need this law for our case we have \[(2^3)^{x+1}\]

OpenStudy (xapproachesinfinity):

what can you do

OpenStudy (anonymous):

One moment ill brb

OpenStudy (xapproachesinfinity):

no problem take your time and reflect

OpenStudy (anonymous):

I am really sorry I have to go, my mother wants us to go out. I thank you for your help though!

OpenStudy (anonymous):

I will try to work on it later and tag you in it

OpenStudy (xapproachesinfinity):

no problem:) anytime!